A box contains 5 balls. Two are numbered 1, one is numbered 2, and two are numbered 3. The balls are mixed and one is selected at random. After a ball is selected, its number is recorded. Then it is replaced. If the experiment is repeated many times, find the mean, variance, and standard deviation of the numbers on the balls.


1
Expert's answer
2021-09-22T00:58:21-0400

Solution:

Let X be the number on each ball. The probability distribution is



The mean is

 "\mu=\Sigma X \cdot P(X)=1 \cdot \frac{2}{5}+2 \cdot \frac{1}{5}+3 \cdot \frac{2}{5}=2"  

The variance is

"\begin{aligned}

\sigma &=\Sigma\left[X^{2} \cdot P(X)\right]-\mu^{2} \\

&=1^{2} \cdot \frac{2}{5}+2^{2} \cdot \frac{1}{5}+3^{2} \cdot \frac{2}{5}-2^{2} \\

&= \frac{24}{5}-4 \\

&=\frac{4}{5}

\end{aligned}"  

The standard deviation is

 "\sigma=\sqrt{\frac{4}{5}}=\sqrt{0.8}=0.894"